I have a parabola that I need to find an equation for.
so far all i have is y=negative___x squared+6. I need to fill in the blank.
I know it passes through (neg.4,5) and (7,2)
help me figure out what goes in the blank please
I just cannot remember how to do it out

Answer by venugopalramana(3286)

I have a parabola that I need to find an equation for.
so far all i have is y=negative___x squared+6. I need to fill in the blank.
I know it passes through (neg.4,5) and (7,2)
help me figure out what goes in the blank please
I just cannot remember how to do it out
Y=-A*X^2+6...IF IT PASSES THROUGH (-4,5)
5= -A *(-4)^2+6 = -16A+6
16A=6-5=1
A=1/16.....
LET US CHECK WHETHER (7,2) LIES ON THAT
2=-(1/16)*7^2+6 IS NOT CORRECT ..SO PLEASE CHECK YOUR DATA
BUT THIS IS THE METHOD OF SOLUTION.SINCE THERE IS ONLY 1 UNKNOWN WE NEED ONLY ONE POINT..THE SECOND POINT IS REDUNDANT..